# Supplemetary Material for the paper

### v1

@@ C.2 Proof of Corollary 1
Assume the conditions of Theorem 5.1, and in addition assume that all reference trajectories $\tilde{\mathbf{X}}(t)$ converge to a solution $\mathbf{X}^{\star}\in\mathbb{R}^{N\times L}$. This convergence is uniform in the sense that the trajectories are confined to a bounded domain, and for any $\epsilon_{app}>0$ there exists a time $T_{c}\leq T$ after which they are all within a distance $\epsilon_{app}$ from $\mathbf{X}^{\star}$. Then for any $\epsilon_{app}>0$, if the initialization scales $\alpha_{V},\alpha_{g}$ are sufficiently small, for any time $t\in[T_{c},T]$ it holds that $\left\|\mathbf{X}(t)-\mathbf{X}^{\star}\right\|_{F}\leq 2\epsilon_{app}$
@@
For $\epsilon_{app}>0$, there exists a time $T_{c}>0$ after which all the reference trajectories are within a distance of $\epsilon_{app}$ from the solution $\mathbf{X}^{\star}$. i.e.,
$\|\tilde{\mathbf{X}}(t)-\mathbf{X}^{\star}\|\leq\epsilon_{app}.$ (257)
@@
$\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|\leq\epsilon_{app}<1.$ (258)
This bound holds atleast until $t\geq T$ or $\|\mathbf{X}(t)\|\geq D+1$. In line with the analysis done by razin2021implicit, by the way of contradiction, we show that the second condition cannot hold, namely $\|\mathbf{X}(t)\|\leq D+1$ cannot be true for any $t\in[0,T]$. Let $t^{\prime}\in[0,T]$ be an initial time at which $\|\mathbf{X}(t^{\prime})\|\geq D+1$. Since $t^{\prime}\leq T$, the guarantee from Theorem 5.1 is still valid, meaning, $\|\mathbf{X}(t^{\prime})-\tilde{\mathbf{X}}(t^{\prime})\|<1$, using the triangle inequality we have:
@@
This is a contradiction to the assumption that $\tilde{\mathbf{X}}(t)$ is confined to a ball of radius $D$ for $t=t^{\prime}$. Thus, for all $t\in[0,T]$, the guarantee $\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|\leq\epsilon_{app}$ holds.
@@
$\|\mathbf{X}(t)-\mathbf{X}^{\star}\|\leq\|\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|+\|\tilde{\mathbf{X}}(t)-\mathbf{X}^{\star}\|\leq 2\epsilon_{app}$ (261)
@@ Appendix D Result on MNIST
To demonstrate the efficacy of the proposed algorithm on real-world data, we conduct a recovery experiment on a set of images from the MNIST dataset. The experiment uses the first $500$ images from the MNIST dataset, each reshaped into a vector of $784$ pixels. The images are normalized to have pixel values between 0 and $1$. A random Gaussian sensing matrix $\mathbf{A}$ of size $1024\times 784$ with $\ell_{2}$-normalized columns is used to generate the measurements. The measurement matrix $\mathbf{Y}$ is then computed as $\mathbf{Y}=\mathbf{A}\mathbf{X}$, where $\mathbf{X}$ represents the matrix of vectorized MNIST images. The proposed IR-MMV algorithm is used to recover the images from the measurements $\mathbf{Y}$ and the sensing matrix $\mathbf{A}$. For comparison, we also show the results from the M-OMP, M-SP and M-FOCUSS algorithm. For M-OMP and M-SP, which require prior knowledge of the row sparsity level, is set to $K=18$. The simulations were performed on the CPU of a desktop computer equipped with an Intel Core i77-$12700$ processor, 1616GB of DDR55 RAM. The MNIST dataset was obtained under the Creative Commons Attribution-Share Alike 3.0 license.

### v2

@@ C.3 Proof of Corollary 1
Assume the conditions of Theorem 5.6, and in addition assume that all reference trajectories $\tilde{\mathbf{X}}(t)$ converge to a solution $\mathbf{X}^{\star}\in\mathbb{R}^{N\times L}$. This convergence is uniform in the sense that the trajectories are confined to a bounded domain, and for any $\epsilon_{\text{app}}>0$ there exists a time $T_{c}\leq T$ after which they are all within a distance $\epsilon_{\text{app}}$ from $\mathbf{X}^{\star}$. Then for any $\epsilon_{\text{app}}>0$, if the initialization scales $\alpha_{V},\alpha_{g}$ are sufficiently small, for any time $t\in[T_{c},T]$ it holds that $\left\|\mathbf{X}(t)-\mathbf{X}^{\star}\right\|_{F}\leq 2\epsilon_{\text{app}}$
@@
For $\epsilon_{\text{app}}>0$, there exists a time $T_{c}>0$ after which all the reference trajectories are within a distance of $\epsilon_{\text{app}}$ from the solution $\mathbf{X}^{\star}$. i.e.,
$\|\tilde{\mathbf{X}}(t)-\mathbf{X}^{\star}\|\leq\epsilon_{\text{app}}.$ (237)
@@
$\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|\leq\epsilon_{\text{app}}<1.$ (238)
This bound holds atleast until $t\geq T$ or $\|\mathbf{X}(t)\|\geq D+1$. In line with the analysis done by Razin et al. (2021), by the way of contradiction, we show that the second condition cannot hold, namely $\|\mathbf{X}(t)\|\leq D+1$ cannot be true for any $t\in[0,T]$. Let $t^{\prime}\in[0,T]$ be an initial time at which $\|\mathbf{X}(t^{\prime})\|\geq D+1$. Since $t^{\prime}\leq T$, the guarantee from Theorem 5.6 is still valid, meaning, $\|\mathbf{X}(t^{\prime})-\tilde{\mathbf{X}}(t^{\prime})\|<1$, using the triangle inequality we have:
@@
This is a contradiction to the assumption that $\tilde{\mathbf{X}}(t)$ is confined to a ball of radius $D$ for $t=t^{\prime}$. Thus, for all $t\in[0,T]$, the guarantee $\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|\leq\epsilon_{\text{app}}$ holds.
@@
$\|\mathbf{X}(t)-\mathbf{X}^{\star}\|\leq\|\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|+\|\tilde{\mathbf{X}}(t)-\mathbf{X}^{\star}\|\leq 2\epsilon_{\text{app}}$ (241)
@@ Appendix D Result on MNIST
To demonstrate the efficacy of the proposed algorithm on real-world data, we conduct a recovery experiment on a set of images from the MNIST dataset. The experiment uses the first $100$ images from the MNIST dataset, each reshaped into a vector of $784$ pixels. The images are normalized to have pixel values between 0 and $1$. A random Gaussian sensing matrix $\mathbf{A}$ of size $700\times 784$ with $\ell_{2}$-normalized columns is used to generate the measurements. The measurement matrix $\mathbf{Y}$ is then computed as $\mathbf{Y}=\mathbf{A}\mathbf{X}$, where $\mathbf{X}$ represents the matrix of vectorized MNIST images. The proposed IR-MMV algorithm is used to recover the images from the measurements $\mathbf{Y}$ and the sensing matrix $\mathbf{A}$. For comparison, we also show the results from the M-OMP, M-SP, M-FOCUSS, AMP-MMV, and M-BSBL algorithms. For M-OMP, M-SP, and AMP-MMV which require prior knowledge of the row sparsity level, it is set to $K=300$. The simulations were performed on the GPU of a desktop computer equipped with an Intel Core i55-$7600K$ processor, 1616GB of DDR44 RAM and GeForce GTX $1070$ with 88GB RAM. The MNIST dataset was obtained under the Creative Commons Attribution-Share Alike 3.0 license.
@@ Appendix E Code and Reproducibility
M-BSBL Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AMP-MMV Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption